Deposits are made in a geometric progression. Given that 4th and the 7th deposits are 24000 and 192000 respectively. Calculate
1- The 10th deposits
2- The sum of the first 10 deposits
1
Expert's answer
2020-12-13T19:21:54-0500

let "a=" first deposit, "r=" common ratio

"4th" deposit correspond to "ar^{3}" and "7th" correspond to "ar^{6}"

"ar^{3}=24000" and "ar^{6}=192000"

"a=24000\/r^{3}" and "a=192000\/r^{6}"

"24000\/r^{3}=192000\/r^{6}"

"r^{3}=8, r=8^{1\/3}=2"

"a=24000\/8=3000"

1) the tenth deposit "=ar^{9} = 3000\times 2^{9}=1,536,000"

2) sum of the first 10 deposits = "a(r^{n}-1)\/(r-1)" "=3000(2^{10}-1)\/(2-1)= 3,069,000"